OpenStudy (janelle):

i want to know what is ment by the perfect square. in equations such as: 4x^2+2(2x)(3b)+9b^2 and x^2+2(x)(4)+4^2

OpenStudy (blacksteel):

A perfect square is a number that is equal to a whole number (or whole number times variable) squared. For example, 64 and 16x^4 are perfect squares because 64 = 8^2 and 16x^4 = (4x^2)^2

OpenStudy (blacksteel):

In the above equations, 4x^2, 9b^2, x^2, and 4^2 are all perfect squares, since when you take the square roots of them you get either an integer or a variable with an integer coefficient.

OpenStudy (blacksteel):

Similarly, perfect cubes, perfect powers of 4, etc are the same thing for the cubic root, 4th root, etc For example, 8 = 2^3 is a perfect cube and 81 = 3^4 is a perfect 4th power

OpenStudy (campbell_st):

a perfect square is a binomal multiplied by itself e.g (2x + 5)(2x5) can be written (2x+5)^2

OpenStudy (blacksteel):

Campbell is referring to perfect square trinomials, which are 3-term polynomials that can be factored into the square of a single binomial.

OpenStudy (blacksteel):

Ex.: 4x^2 + 12x + 9 = (2x + 3)^2

OpenStudy (janelle):

can you complete this square in detail for me plz? (x+h)^2+R and explain the minimum and maximum value thing?

OpenStudy (blacksteel):

Well, (x+h)^2 = (x+h)(x+h) To do this, you multiply each pair of terms together and add them all up. So you get x*x + x*h + h*x + h*h = x^2 + 2xh + h^2 So (x+h)^2 + R = x^2 + h^2 + 2xh + R I'm not sure what you mean by minimum and maximum value - can you give me more info? (I also have to drive home soon, so I won't be able to do any more answers right away)

OpenStudy (janelle):